# Scale-Dependent Radial Geometry and Metric Mismatch in Wasserstein Propagation for Reverse Diffusion

### 20 Mar 2026 v1

@@ Appendix D Switch-time conversion back to $W_{2}$
D.1 One explicit admissible moment budget
@@ Appendix E Late-window Euclidean propagation
Fix $s_{0}\in[0,T]$, let $t_{s}:=T-s_{0}$, and assume Assumptions 1, 2, and 3. Assume moreover that $\Gamma(s_{0})<\infty$, that the learned reverse SDE is strongly well posed on $t_{s},T]$, that the theorem-level ideal reverse SDE (12) admits a global strong solution
@@ F.1 Proof of the main theorem and switch selection
Corollary 2 (Discrete one-switch selection principle).
@@
Consequently, whenever $\mathcal{S}_{\mathrm{adm}}^{h}\neq\varnothing$, there exists $s_{0}^{\star}\in\mathcal{S}_{\mathrm{adm}}^{h}$ such that
@@
Remark 8 (Why one-switch selection is nontrivial).
Admissibility and optimization play different roles. Admissibility is a geometric existence statement: it says that on the whole early window the far field carries a positive contraction reserve. Optimization is a geometric routing statement within the one-switch family studied in this paper: it decides when one should stop using the soft radial metric, pay the one-time conversion back to $W_{2}$, and expose the remaining error to the late Euclidean load. We do not claim here that this scalar optimization is globally optimal among multi-switch or continuously varying metric strategies.
@@
The first claim is immediate from Theorem 1 and Definition 7. Since $\mathcal{S}_{\mathrm{adm}}^{h}$ is finite whenever nonempty, $\mathcal{B}_{p}$ attains a minimum on $\mathcal{S}_{\mathrm{adm}}^{h}$, and applying Theorem 1 at a minimizer gives the second claim. ∎
F.2 Formal comparison with the direct full-horizon Euclidean route
Proposition 5 (Direct full-horizon Euclidean propagation).
@@ F.2 Formal comparison with the direct full-horizon Euclidean route
Corollary 3 (Formal comparison with the direct full-horizon Euclidean route).
Assume the hypotheses of Theorem 1 and of Proposition 5. Fix an admissible grid-aligned switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, with $t_{K}=t_{s}:=T-s_{0}$. Define
$\mathcal{D}_{\mathrm{dir}}(s_{0})$={} $\Delta_{\mathrm{late}}^{\mathrm{disc}}(s_{0})+\Delta_{\mathrm{late}}^{\mathrm{force}}(s_{0})$ (49)
$+e^{\Gamma(T)}W_{2}(\widehat{p}_{T},p_{T})+\sum_{k=0}^{K-1}e^{\Gamma(T-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}+\int_{s_{0}}^{T}e^{\Gamma(s)}\,g^{2}(s)\,\varepsilon(s)\,\mathop{}\!\mathrm{d}s.$
@@
The phase-aware part is exactly Theorem 1 packaged by Definition 7. The direct part comes from Proposition 5 by splitting the sum and the integral at the grid index $K$. ∎
Remark 9 (What the comparison isolates).
The direct Euclidean route and the phase-aware route differ only in how they transport the pre-switch terms. The direct route sends initialization mismatch, pre-switch discretization, and pre-switch score forcing directly through the full-horizon Euclidean weights. By contrast, the phase-aware route first compresses the same three inputs inside $\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0})$, pays the one-time conversion back to $W_{2}$, and only then incurs the shorter late-window factor $e^{\Gamma(s_{0})}$. Thus the gain is structural: it comes from routing the same early errors through different geometries.
Appendix G Operational supplements: proxy switches and verification
Nothing in this section is needed for the proof of Theorem 1. Its role is operational: to compare Assumption 3 with more classical global score-regularity hypotheses, to record conservative proxy switches, to give verifiable sufficient conditions for the early reflection-coupling input, and to exhibit a concrete worked example in which the one-switch theorem can be instantiated explicitly.
G.1 Assumption 3 compared with global score regularity
Remark 10 (Structural versus verification assumptions).
Assumption 3 is the structural geometric hypothesis used by the theorem. It enters only through the one-sided loss $\ell(s)$, which perturbs the radial profile and the late Euclidean load. By contrast, global Lipschitz assumptions on the score error $e_{s}$, or on the learned score $s_{\theta}(\cdot,s)$ together with the true score $\nabla\log p_{s}$, are stronger verification devices. They are useful because they imply Assumption 3 and also fit naturally with the well-posedness and reflection-coupling input used later, but they are not the quantities that the geometric argument itself sees.
Lemma 16 (Jacobian and Lipschitz criteria for Assumption 3).
@@ G.1 Assumption 3 compared with global score regularity
    $\lambda_{\max}\!\Big(\tfrac{1}{2}\big(\nabla e_{s}(x)+\nabla e_{s}(x)^{\top}\big)\Big)\leq\bar{\ell}(s),\qquad x\in\mathbb{R}^{d};$
@@
    there exist measurable $L_{\theta},L_{\star}:(0,T]\to[0,\infty)$ such that for $x,y\in\mathbb{R}^{d}$
    $\|s_{\theta}(x,s)-s_{\theta}(y,s)\|\leq L_{\theta}(s)\|x-y\|,\qquad\|\nabla\log p_{s}(x)-\nabla\log p_{s}(y)\|\leq L_{\star}(s)\|x-y\|.$
@@
G.2 A computable proxy for admissibility and switch selection
Proposition 6 (Conservative proxy switches and proxy objective).
@@ G.2 A computable proxy for admissibility and switch selection
G.3 Verifiable sufficient conditions for the early reflection-coupling input
@@
Corollary 4 (A network-level checklist).
@@
G.4 Worked example: a variance-preserving schedule with Euler/Heun-type defects
Proposition 8 (Explicit proxy geometry and a concrete admissible regime in the VP example).
@@ G.4 Worked example: a variance-preserving schedule with Euler/Heun-type defects
for constants $\bar{\ell}\in\mathbb{R}$ and $\bar{\varepsilon}\geq 0$. Define
@@
Then
$a(s)=e^{-\beta s/2},\qquad\sigma^{2}(s)=1-e^{-\beta s},\qquad\alpha_{s}=\frac{\alpha}{D(s)},\qquad M_{s}=\frac{Me^{-\beta s}}{D(s)^{2}}.$
Consequently the proxy reserve and load are
@@
If $\alpha<1$ and $\bar{\ell}<\frac{1}{2}$, define
@@
Then every grid-aligned $s_{0}\in[s_{\mathrm{pr}}^{\star},T]$ is proxy-admissible, provided $s_{\mathrm{pr}}^{\star}\leq T$.
Moreover, the proxy late-window load is explicit:
@@
If, in addition, the numerical grid is uniform, $t_{k}=kh$, and
$\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}\leq C_{\mathrm{sch}}h^{q},\qquad k=0,\dots,N-1,$
then every proxy-admissible grid-aligned switch $s_{0}$ yields a concrete one-switch bound
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})+e^{\Gamma_{\mathrm{pr}}(s_{0})}\,C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})\,\Bigl(\overline{\Delta}_{\varphi,\mathrm{pr}}^{\mathrm{sw}}(s_{0})\Bigr)^{\theta_{p}},$
with
$\overline{\Delta}_{\varphi,\mathrm{pr}}^{\mathrm{sw}}(s_{0})\lesssim e^{-c_{\mathrm{pr}}(s_{0})(T-s_{0})}W_{\varphi_{s_{0}}^{\mathrm{pr}}}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}(T-s_{0})h^{q-1}+\frac{\beta\bar{\varepsilon}}{c_{\mathrm{pr}}(s_{0})}\Bigl(1-e^{-c_{\mathrm{pr}}(s_{0})(T-s_{0})}\Bigr),$
and
$\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})\lesssim C_{\mathrm{sch}}\,s_{0}\,h^{q-1}\,\sup_{u\in[0,s_{0}]}e^{\Gamma_{\mathrm{pr}}(u)}+\beta\bar{\varepsilon}\int_{0}^{s_{0}}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u.$
By contrast, the direct full-horizon Euclidean route gives only
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\lesssim e^{\Gamma_{\mathrm{pr}}(T)}W_{2}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}\,T\,h^{q-1}\,\sup_{u\in[0,T]}e^{\Gamma_{\mathrm{pr}}(u)}+\beta\bar{\varepsilon}\int_{0}^{T}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u.$
@@
Substituting these into (16) gives
$\alpha_{s}=\frac{\alpha}{\alpha+(1-\alpha)e^{-\beta s}}=\frac{\alpha}{D(s)},\qquad M_{s}=\frac{Me^{-\beta s}}{D(s)^{2}}.$
The formulas for $\mathfrak{m}_{\mathrm{pr}}$ and $\mathfrak{b}_{\mathrm{pr}}$ follow immediately from Definition 3 with $\ell$ replaced by $\bar{\ell}$.
Assume now $\alpha<1$. Then $s\mapsto\alpha_{s}$ is increasing, so
@@
Solving this inequality gives exactly $s_{0}\geq s_{\mathrm{pr}}^{\star}$.
For the load, write

### 30 Mar 2026 v2

@@ Appendix D Switch-time conversion back to $W_{2}$
D.1 An explicit admissible moment criterion
@@ Appendix E Late-window Euclidean propagation
Fix $s_{0}\in[0,T]$ and assume that the grid contains the switch time
$t_{s}:=T-s_{0}=t_{K}$
for some $K\in\{0,\dots,N\}$. Assume Assumptions 1, 2, and 3. Assume moreover that $\Gamma(s_{0})<\infty$, that the learned reverse SDE is strongly well posed on $t_{s},T]$, that the ideal reverse SDE (12) admits a global strong solution
@@ F.1 Proof of the main theorem and switch selection
Corollary 2 (Discrete one-switch routing principle).
@@
Hence, whenever $\mathcal{S}_{\mathrm{adm}}^{h}\neq\varnothing$, the one-switch routing problem on the numerical grid reduces to the scalar minimization
$\min_{s_{0}\in\mathcal{S}_{\mathrm{adm}}^{h}}\mathcal{B}_{p}(s_{0}).$
In particular, there exists $s_{0}^{\star}\in\mathcal{S}_{\mathrm{adm}}^{h}$ such that
@@
Remark 8.
The substance of Corollary 2 is not the elementary fact that a minimum exists on a finite set. Its content is structural: Theorem 1 has already compressed an end-to-end geometric routing problem into the single scalar field
[…]
Admissibility is the geometric existence statement that the early window supports a positive far-field contraction reserve. Optimization is the quantitative routing statement, within the one-switch family studied here, that chooses where one should leave the first contracting geometry, pay the one-time conversion back to $W_{2}$, and expose the remaining discrepancy to the late Euclidean load. In this sense, the corollary turns a pathwise choice of interface between two geometries into a one-dimensional design problem on the admissible grid.
It also clarifies why an exact routed-versus-direct comparison is possible: for fixed $s_{0}$, the late-window contribution is the same in both bounds, so any strict improvement can only come from the early window. We do not claim that this scalar optimization is globally optimal among multi-switch or continuously varying metric strategies.
@@
Theorem 1 together with Definition 7 yields
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\mathcal{B}_{p}(s_{0}),\qquad s_{0}\in\mathcal{S}_{\mathrm{adm}}^{h}.$
If $\mathcal{S}_{\mathrm{adm}}^{h}\neq\varnothing$, then $\mathcal{S}_{\mathrm{adm}}^{h}$ is finite, hence $\mathcal{B}_{p}$ attains a minimum on $\mathcal{S}_{\mathrm{adm}}^{h}$. Applying Theorem 1 at a minimizer gives the stated bound. ∎
F.2 Comparison with direct Euclidean propagation
Proposition 5.
Assume moreover that the ideal reverse SDE (12) admits a global strong solution $(\widetilde{X}_{t})_{0\leq t\leq T}$ with
$\operatorname{Law}(\widetilde{X}_{t})=p_{T-t},\qquad 0\leq t\leq T.$
@@ F.2 Comparison with direct Euclidean propagation
Corollary 3.
Assume the hypotheses of Theorem 1 and of Proposition 5. Fix an admissible grid-aligned switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, with $t_{K}=t_{s}:=T-s_{0}$. Write
$d_{k}:=\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2},\qquad\Lambda_{s_{0}}(r):=\int_{s_{0}}^{r}\mathfrak{b}(u)\,\mathop{}\!\mathrm{d}u,\qquad r\in[s_{0},T].$
Define the common late-window term
$\mathcal{L}(s_{0}):=\Delta_{\mathrm{late}}^{\mathrm{disc}}(s_{0})+\Delta_{\mathrm{late}}^{\mathrm{force}}(s_{0}),$
the routed early transport term
$\mathcal{R}_{\mathrm{sw}}(s_{0}):=C_{p}^{\mathrm{sw}}(s_{0})\Biggl[e^{-c(s_{0})t_{s}}W_{\varphi_{s_{0}}}(\widehat{p}_{T},p_{T})+\sum_{k=0}^{K-1}e^{-c(s_{0})(t_{s}-t_{k+1})}d_{k}+\int_{s_{0}}^{T}e^{-c(s_{0})(s-s_{0})}g^{2}(s)\varepsilon(s)\,\mathop{}\!\mathrm{d}s\Biggr]^{\theta_{p}},$
and the direct early transport term
$\mathcal{R}_{\mathrm{dir}}(s_{0}):=e^{\Lambda_{s_{0}}(T)}W_{2}(\widehat{p}_{T},p_{T})+\sum_{k=0}^{K-1}e^{\Lambda_{s_{0}}(T-t_{k+1})}d_{k}+\int_{s_{0}}^{T}e^{\Lambda_{s_{0}}(s)}g^{2}(s)\varepsilon(s)\,\mathop{}\!\mathrm{d}s.$
@@
$\mathcal{B}_{p}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{sw}}(s_{0}),\qquad\mathcal{D}_{\mathrm{dir}}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{dir}}(s_{0}),$
and hence
@@
Moreover,
$\mathcal{B}_{p}(s_{0})\leq\mathcal{D}_{\mathrm{dir}}(s_{0})\iff\mathcal{R}_{\mathrm{sw}}(s_{0})\leq\mathcal{R}_{\mathrm{dir}}(s_{0}),$
with strict inequality if and only if $\mathcal{R}_{\mathrm{sw}}(s_{0})<\mathcal{R}_{\mathrm{dir}}(s_{0})$.
@@
The identity
$\mathcal{B}_{p}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{sw}}(s_{0})$
is just Definition 7. For the direct route, note that
$e^{\Gamma(T)}=e^{\Gamma(s_{0})}e^{\Lambda_{s_{0}}(T)},\qquad e^{\Gamma(T-t_{k+1})}=e^{\Gamma(s_{0})}e^{\Lambda_{s_{0}}(T-t_{k+1})},$
and
$e^{\Gamma(s)}=e^{\Gamma(s_{0})}e^{\Lambda_{s_{0}}(s)},\qquad s\in[s_{0},T].$
Substituting these identities gives
$\mathcal{D}_{\mathrm{dir}}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{dir}}(s_{0}).$
The bound by $\min\{\mathcal{B}_{p}(s_{0}),\mathcal{D}_{\mathrm{dir}}(s_{0})\}$ combines Theorem 1 with Proposition 5, and the comparison criterion follows by subtracting the common term $\mathcal{L}(s_{0})$ and then the common factor $e^{\Gamma(s_{0})}$. ∎
Remark 9.
The comparison becomes transparent only after factoring out the common post-switch factor $e^{\Gamma(s_{0})}$. After that reduction, the two routes propagate the same three pre-switch inputs in opposite geometric directions. The direct route weights an input created at noise level $u\in[s_{0},T]$ by the residual Euclidean factor
$e^{\Lambda_{s_{0}}(u)},$
whereas the routed bound first damps the same input by
$e^{-c(s_{0})(u-s_{0})},$
and then pays a single return to $W_{2}$ through $C_{p}^{\mathrm{sw}}(s_{0})$ and $\theta_{p}$.
This identifies the genuine mismatch regime: radial contraction is already available, while direct Euclidean propagation on the same window is still net expansive. A convenient sufficient indicator is
$\underline{m}(s_{0})>0\qquad\text{and}\qquad\Lambda_{s_{0}}(T)=\int_{s_{0}}^{T}\mathfrak{b}(v)\,\mathop{}\!\mathrm{d}v>0.$
Then the routed bound damps the early inputs, whereas the direct bound amplifies them on net. The only competing loss is the one-time metric conversion. Accordingly, a strict improvement should be expected when the early window is long enough and sufficiently inside the admissible regime so that $c(s_{0})$ is appreciable and $C_{p}^{\mathrm{sw}}(s_{0})$ is not too large. By contrast, if the residual Euclidean load is already nonpositive, or if $s_{0}$ is chosen too close to the admissibility threshold, there is no reason to expect a systematic advantage from routing. This regime is made explicit in the variance-preserving example of Proposition˜8.
Appendix G Auxiliary comparison principles and sufficient conditions
Nothing in this section is needed for the proof of Theorem 1. It records auxiliary comparison principles and sufficient conditions: a relation between Assumption 3 and classical global score-regularity hypotheses, conservative admissibility certificates for switch selection, sufficient conditions for the early reflection-coupling hypothesis, and an explicit variance-preserving comparison result.
G.1 Relation of score-error monotonicity to global score regularity
Remark 10 (What Assumption 3 actually controls).
Assumption 3 is the theorem-level geometric input on the score error. Although it is stated in one-sided Lipschitz form, what the argument actually uses is more specific: for a pair $(x,y)$, the score-error contribution to pairwise distance evolution appears only through
$g^{2}(s)\,\langle e_{s}(x)-e_{s}(y),x-y\rangle.$
Accordingly, the radial lower profile $\underline{\kappa}_{s}$, the switch reserve $\mathfrak{m}(s)$, and the late Euclidean load $\mathfrak{b}(s)$ depend on $e_{s}$ only through the component of $e_{s}(x)-e_{s}(y)$ parallel to the separation vector xyx-y. Purely transverse or rotational components are invisible to this mechanism. In this sense Assumption 3 is a pairwise geometric monotonicity condition, closer to the one-sided dissipativity hypotheses standard in SDE analysis [Higham et al., 2002] than to a full global Lipschitz bound on the vector field. By contrast, global Lipschitz assumptions are stronger verification devices, and are also the natural hypotheses under which direct Euclidean $W_{2}$ arguments are typically closed [Gao et al., 2025, Gao and Zhu, 2024].
Lemma 16 (Sufficient condition for Assumption 3).
@@ G.1 Relation of score-error monotonicity to global score regularity
    $\lambda_{\max}\!\Bigl(\tfrac{1}{2}\bigl(\nabla e_{s}(x)+\nabla e_{s}(x)^{\top}\bigr)\Bigr)\leq\bar{\ell}(s),\qquad x\in\mathbb{R}^{d};$
@@
    there exist measurable $L_{\theta},L_{\star}:(0,T]\to[0,\infty)$ such that
    $\|s_{\theta}(x,s)-s_{\theta}(y,s)\|\leq L_{\theta}(s)\|x-y\|,\qquad\|\nabla\log p_{s}(x)-\nabla\log p_{s}(y)\|\leq L_{\star}(s)\|x-y\|,\qquad x,y\in\mathbb{R}^{d}.$
@@
In particular, the symmetrized Jacobian bound, global Lipschitz control of $e_{s}$, and separate Lipschitz control of $s_{\theta}$ and $\nabla\log p_{s}$ all imply Assumption 3, but they are not necessary for it.
@@
Remark 11 (Assumption 3 is strictly weaker than global Lipschitz).
The converse of Lemma 16 fails dramatically. Let
$e_{s}(x)=A_{s}x,\qquad A_{s}^{\top}=-A_{s},$
where $\|A_{s}\|_{\mathrm{op}}$ may be arbitrarily large. Then for every $x,y\in\mathbb{R}^{d}$,
$\langle e_{s}(x)-e_{s}(y),x-y\rangle=\langle A_{s}(x-y),x-y\rangle=0,$
so Assumption 3 holds with $\ell(s)=0$. However,
$\|e_{s}(x)-e_{s}(y)\|\leq\|A_{s}\|_{\mathrm{op}}\|x-y\|,$
and the global Lipschitz constant can be arbitrarily large. Thus Assumption 3 controls only the symmetric/expansive part of the score error, not the full vector increment. This is exactly why it is the right structural condition for pairwise distance propagation.
Remark 12 (Assumptions 2 and 3 are complementary).
Assumption 2 and Assumption 3 address different error channels. The former is an averaged forcing bound under $p_{s}$:
$\mathop{\mathbb{E}}\nolimits_{X\sim p_{s}}\|e_{s}(X)\|^{2}\leq\varepsilon^{2}(s).$
The latter is a pairwise geometric monotonicity bound:
$\langle e_{s}(x)-e_{s}(y),x-y\rangle\leq\ell(s)\|x-y\|^{2}.$
Neither replaces the other. A field may satisfy Assumption 3 with $\ell(s)=0$ and still have large $L^{2}(p_{s})$-size; the skew-symmetric example above already shows this as soon as $p_{s}$ has nontrivial second moment. Conversely, a field may have small averaged $L^{2}(p_{s})$-error while having a large positive one-sided slope on a region of small $p_{s}$-mass; such a field barely changes the forcing integral but can destroy pairwise contraction. So the two hypotheses are not redundant: Assumption 2 controls averaged forcing, while Assumption 3 controls pairwise expansion.
G.2 Conservative admissibility and switch-selection certificates
Proposition 6 (Proxy admissibility and proxy certificates).
@@ G.2 Conservative admissibility and switch-selection certificates
G.3 Sufficient conditions for the early reflection-coupling hypothesis
@@
Corollary 4 (A score-model sufficient condition).
@@
G.4 Variance-preserving schedules: explicit routed-versus-direct certificates
The point of this example is not merely to compute proxy admissibility explicitly. It is to exhibit a concrete regime in which the phase-aware certificate is structurally sharper than the direct Euclidean certificate. For the variance-preserving schedule, the formulas become fully explicit. In the regime $\alpha<1$, Gaussian smoothing increases $\alpha_{s}$ monotonically, so both the opening of the admissible radial window and the routed-versus-direct comparison can be written in closed form.
Proposition 8 (VP proxy formulas and explicit routed-versus-direct certificates).
@@ G.4 Variance-preserving schedules: explicit routed-versus-direct certificates
for constants $\bar{\ell}\in\mathbb{R}$ and $\bar{\varepsilon}\geq 0$. Let
@@
Assume moreover that the grid is uniform, $t_{k}=kh$, and that
$\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}\leq C_{\mathrm{sch}}h^{q},\qquad k=0,\dots,N-1.$
For a proxy-admissible grid-aligned switch $s_{0}$, let $t_{s}:=T-s_{0}=t_{K}$, $L:=T-s_{0}$, and let $\varphi_{s_{0}}^{\mathrm{pr}}$, $c_{\mathrm{pr}}(s_{0})$, and $C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})$ be the proxy switch metric, damping rate, and conversion constant furnished by Proposition 6 and (34)–(35) after replacing
Then the following hold.
  1. (i)
    The smoothed weak-log-concavity parameters are
    $a(s)=e^{-\beta s/2},\qquad\sigma^{2}(s)=1-e^{-\beta s},\qquad\alpha_{s}=\frac{\alpha}{D(s)},\qquad M_{s}=\frac{Me^{-\beta s}}{D(s)^{2}},$
    and the proxy reserve and load are
@@
  2. (ii)
  3. (iii)
    Suppose $\alpha<1$ and $\bar{\ell}<\frac{1}{2}$. Define
@@
    Then every grid-aligned $s_{0}\in[s_{\mathrm{pr}}^{\star},T]$ is proxy-admissible, provided $s_{\mathrm{pr}}^{\star}\leq T$. Equivalently, proxy admissibility begins exactly when
    $\alpha_{s_{0}}>\bar{\ell}+\frac{1}{2}.$
    For later use, it is also convenient to record the primitive
@@
    The second branch is included only to record the closed-form primitive in the degenerate case $\alpha=1$.
  4. (iv)
    Define the common late-window proxy term
    $\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0}):=C_{\mathrm{sch}}h^{q}\sum_{k=K}^{N-1}e^{\Gamma_{\mathrm{pr}}(T-t_{k+1})}+\beta\bar{\varepsilon}\int_{0}^{s_{0}}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u,$
    and the routed early budget
    $\Xi_{\mathrm{sw}}^{\mathrm{pr}}(s_{0}):=e^{-c_{\mathrm{pr}}(s_{0})L}\,W_{\varphi_{s_{0}}^{\mathrm{pr}}}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}h^{q}\,\frac{1-e^{-c_{\mathrm{pr}}(s_{0})L}}{1-e^{-c_{\mathrm{pr}}(s_{0})h}}+\frac{\beta\bar{\varepsilon}}{c_{\mathrm{pr}}(s_{0})}\Bigl(1-e^{-c_{\mathrm{pr}}(s_{0})L}\Bigr).$
    Then the phase-aware proxy certificate satisfies
    $W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\mathcal{B}_{p}^{\mathrm{pr}}(s_{0})\leq\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})+e^{\Gamma_{\mathrm{pr}}(s_{0})}\,C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})\,\Bigl(\Xi_{\mathrm{sw}}^{\mathrm{pr}}(s_{0})\Bigr)^{\theta_{p}}.$
  5. (v)
    Define the direct Euclidean proxy certificate by
    $\mathcal{D}_{\mathrm{dir,pr}}(s_{0}):=\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})+e^{\Gamma_{\mathrm{pr}}(T)}W_{2}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}h^{q}\sum_{k=0}^{K-1}e^{\Gamma_{\mathrm{pr}}(T-t_{k+1})}+\beta\bar{\varepsilon}\int_{s_{0}}^{T}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u.$
    If
    $\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0}):=\inf_{u\in[s_{0},T]}\mathfrak{b}_{\mathrm{pr}}(u)>0,$
    then
    $\mathcal{D}_{\mathrm{dir,pr}}(s_{0})-\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}\,\Xi_{\mathrm{dir}}^{\mathrm{pr}}(s_{0}),$
    where
    $\Xi_{\mathrm{dir}}^{\mathrm{pr}}(s_{0}):=e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}\,W_{2}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}h^{q}\,\frac{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1}{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})h}-1}+\frac{\beta\bar{\varepsilon}}{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})}\Bigl(e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1\Bigr).$
    Consequently, whenever
    $C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})\,\Bigl(\Xi_{\mathrm{sw}}^{\mathrm{pr}}(s_{0})\Bigr)^{\theta_{p}}<\Xi_{\mathrm{dir}}^{\mathrm{pr}}(s_{0}),$
    one has the strict certificate-level improvement
    $\mathcal{B}_{p}^{\mathrm{pr}}(s_{0})<\mathcal{D}_{\mathrm{dir,pr}}(s_{0}).$
@@
Substituting into (16) gives
$\alpha_{s}=\frac{\alpha}{\alpha+(1-\alpha)e^{-\beta s}}=\frac{\alpha}{D(s)},\qquad M_{s}=\frac{Me^{-\beta s}}{D(s)^{2}},$
which proves the formulas in (i). The expressions for $\mathfrak{m}_{\mathrm{pr}}$ and $\mathfrak{b}_{\mathrm{pr}}$ are the specialization of Proposition 6 to $f\equiv\beta/2$, $g^{2}\equiv\beta$, and the constant proxy bound $\bar{\ell}$.
Assume now $\alpha<1$. Then $s\mapsto\alpha_{s}$ is increasing, hence so is $s\mapsto\mathfrak{m}_{\mathrm{pr}}(s)$. Therefore
@@
Solving this inequality gives exactly $s_{0}\geq s_{\mathrm{pr}}^{\star}$, proving the admissibility-threshold claim in (ii).
Independently of that monotonicity argument, the primitive $\Gamma_{\mathrm{pr}}$ is obtained from